A student gets an 86% on one test and 73% on the second test. What does the student need to get on the third exam to stay at a B (83%)?

Found 3 solutions by Mathtut, user_dude2008, stanbon:
Answer by Mathtut(3670)

(.86+.73+x)/3=.83
:
1.59+x=2.49
:
x=.90
:
.9(100)=90% is needed

Answer by user_dude2008(1861)

A student gets an 86% on one test and 73% on the second test. What does the student need to get on the third exam to stay at a B (83%)?


(86+73+x)/3>=83

(159+x)/3>=83

159+x>=83(3)

159+x>=249

x>=249-159

x>=90


Ans: Student needs 90% or better.

Answer by stanbon(57246)

A student gets an 86% on one test and 73% on the second test. What does the student need to get on the third exam to stay at a B (83%)?
------------------
Let 3rd exam score be x%.
Equation:
(83 + 73 + x)/3 = 83
156 + x = 249
x = 93
---------------
The student will need 93% on the 3rd exam.
------------------
Cheers,
Stan H.